Can someone show me how to calculate this please. With steps
4. The movie The Godfather grossed $135 million in 1972. The movie Spiderman IIT grossed $336 million in 2007, Assuming a 4% rate of inflation on average, which one earned more in todays dollars?
Answer #1

We need to calculate the future value of these amounts with respect to their release years.

We need to use the following formula:

Future value = present value * (1+r)n

In which r is the rate of interest (here, r = 4% = 0.04) and n is the number of years.

The current year is 2019; so for The Godfather, 'n' = 2019 - 1972 = 47

So, the future value of the gross earning of Godfather =

Present value * (1+r)n = 135 million * (1+0.04)47 = 135 million * (1.04)47 = 135 million * 6.32 = 853.2 million.

For Spiderman III, 'n' = 2019 - 2007 = 12

So, the future value of the gross earning of Godfather =

Present value * (1+r)n = 336 million * (1+0.04)12 = 336 million * (1.04)12 = 336 million * 1.60 = 537.6 million.

Therefore, The Godfather earned more in today's dollar.